### v2

# Confluent Vandermonde matrix and related topics

Abstract
In this note, we explore the connections between the confluent Vandermonde matrix and several mathematical topics, including interpolating polynomials, Hasse derivatives, LU factorization, companion matrices and their Jordan forms, and the Chinese remainder theorem. Using a unified approach based on polynomial evaluations and derivative computations at selected points, we provide accessible proofs that not only clarify key results but also offer insights for both experienced researchers and those new to the subject.
Keywords: Confluent Vandermonde matrix, interpolating polynomials, Hasse derivatives, LU factorization, companion matrix, Jordan form, Chinese remainder theorem.
AMS subject classifications: 15A20, 15A21, 15A23, 33C45
1 Introduction
Let $\Lambda=\{\lambda_{1},\ldots,\lambda_{d}\}$ be a set of $d$ distinct real numbers. The corresponding Vandermonde matrix is the $d\times d$ matrix
$V_{\Lambda}=\begin{bmatrix}1&\lambda_{1}&\cdots&\lambda_{1}^{d-1}\\ 1&\lambda_{2}&\cdots&\lambda_{2}^{d-1}\\ \vdots&\vdots&~{}&\vdots\\ 1&\lambda_{d}&\cdots&\lambda_{d}^{d-1}\\ \end{bmatrix}.$ (1)
The Vandermonde matrix has been used in polynomial interpolation, coding theory, signal processing, differential equations, and recursive relations; see [20, 14, 5, 2, 16] and the references therein. It was also used to give alternative proofs for some classical results. For example, it was used to prove a theorem of Abel [11], whose classical proof is based on the residue theorem. There are many elegant results on the Vandermonde matrix, and one of them is the following formula for the determinant of $\det(V_{\Lambda})$ which appeared in many textbooks
$\det(V_{\Lambda})=\prod_{i<j}(\lambda_{j}-\lambda_{i}).$
Rushanan[20] gave the LU factorization of $V_{\Lambda}$. Many formulas for $V_{\Lambda}^{-1}$ were studied and rediscovered by researchers in pure and applied subjects; see, e.g., [17, 15]. In particular, Parker[19] pointed out that the columns of $V^{-1}$ are exactly the coefficients of the Lagrange polynomials. The generalization of $V_{\Lambda}$ whose columns are assigned with nonconsecutive powers was considered and revealed a strong connection to symmetric functions through Schur functions; see, e.g., [18, 12, 10, 7, 3, 6].
Brand[1] pointed out an interesting relation that $V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}$ is a diagonal matrix, where $C$ is the companion matrix of $(x-\lambda_{1})\cdots(x-\lambda_{d})$. In the same paper, Brand mentioned that if $\lambda_{1},\dots,\lambda_{d}$ are not distinct, one can extend the definition of the Vandermonde matrix to define the confluent Vandermonde matrix $V_{\Lambda}$ using the higher derivatives of the polynomial $(x-\lambda_{1})\cdots(x-\lambda_{d})$ at the points $\lambda_{j}$, which appear more than once; see the definition in Section 2. Then $V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}$ will be the Jordan canonical form of $C_{\Lambda}$.
Researchers have obtained results on the confluent Vandermonde matrix because of its connections and applications to many different subjects. However, these results are scattered across different areas of the mathematical literature. In this paper, we give a systematic introduction to the confluent Vandermonde matrix. In particular, Using a unified approach based on polynomial evaluations and derivative computations at selected points, we provide accessible proofs that not only clarify key results but also offer insights for both experienced researchers and those new to the subject.
2 Evaluations of a polynomial at specific points
Let
$p(x)=c_{0}+c_{1}x+\cdots+c_{d-1}x^{d-1}$ (2)
be a polynomial over $\mathbb{R}$. Let $\Lambda=\{\lambda_{1},\ldots,\lambda_{d}\}$ be a set of $d$ distinct real numbers and $V_{\Lambda}$ be as in Equation 1. Then
$V_{\Lambda}\begin{bmatrix}c_{0}\\ \vdots\\ c_{d-1}\end{bmatrix}=\begin{bmatrix}p(\lambda_{1})\\ \vdots\\ p(\lambda_{d})\end{bmatrix}.$ (3)
Therefore, $V_{\Lambda}$ can be viewed as a transformation of a polynomial of degree at most $d-1$ to its evaluations at the points in $\Lambda$. In fact, $V_{\Lambda}$ is a linear transformation from the linear space of polynomials with degrees bounded by $d-1$ to ${\mathbb{R}}^{d}$. Since $\det(V_{\Lambda})\neq 0$, $V_{\Lambda}$ is invertible and any assignment of $p(\lambda_{1}),\ldots,p(\lambda_{d})$ defines a unique polynomial $p(x)$ of degree bounded by $d-1$.
When $\Lambda$ contains repeated values, the repeated $\lambda_{i}$ should provide not just the information $p(\lambda_{i})$. Instead, more information should be revealed at the point of $\lambda_{i}$, and the evaluations at its derivatives $p(\lambda_{i})$, $p^{\prime}(\lambda_{i})$, $\ldots$, are natural choices. In fact, one may view the higher derivatives at $\lambda_{i}$ as the information obtained by the limit process when several distinct points converge to $\lambda_{i}$.
Note that if $p(x)=c_{0}+c_{1}x+\cdots+c_{d-1}x^{d-1}$, then
$\begin{array}[]{lcrrrr}\frac{1}{0!}\cdot p(x)&=&c_{0}+&c_{1}x+&c_{2}x^{2}+&\cdots+c_{d-1}x^{d-1},\\ \frac{1}{1!}\cdot\frac{d}{dx}p(x)&=&&c_{1}+&c_{2}2x+&\cdots+c_{d-1}(d-1)x^{d-2},\\ \frac{1}{2!}\cdot\frac{d^{2}}{dx^{2}}p(x)&=&&&c_{2}+&\cdots+c_{d-1}\left(\frac{(d-1)(d-2)}{2}\right)x^{d-3},\\ &\vdots&&&&\end{array},$
and, in general,
$\frac{1}{r!}\cdot\frac{d^{r}}{dx^{r}}p(x)=\sum_{i=r}^{d-1}c_{i}\binom{i}{r}x^{i-r}.$
Let $\mathcal{D}^{(r)}=\frac{1}{r!}\cdot\frac{d^{r}}{dx^{r}}$. Then for any $m\leq d$,
$\begin{bmatrix}1&\lambda&\lambda^{2}&~{}&\cdots&\lambda^{d-1}\\ 0&\binom{1}{1}&\binom{2}{1}\lambda&~{}&\cdots&\binom{d}{1}\lambda^{d-2}\\ 0&0&\binom{2}{2}&~{}&\cdots&\binom{d}{2}\lambda^{d-3}\\ ~{}&~{}&~{}&~{}&\vdots&~{}\\ 0&\cdots&0&\binom{m-1}{m-1}&\cdots&\binom{d}{m-1}\lambda^{d-m}\end{bmatrix}\begin{bmatrix}c_{0}\\ \vdots\\ c_{d-1}\end{bmatrix}=\begin{bmatrix}\mathcal{D}^{(0)}p(\lambda)\\ \vdots\\ \mathcal{D}^{(m-1)}p(\lambda)\end{bmatrix}.$ (4)
Definition 2.1.
Let $\lambda$ be a real number and $m,d$ integers with $m\leq d$. Define the $m\times d$ matrix
$V_{\lambda^{(m)}}^{(d)}=\begin{bmatrix}1&\lambda&\lambda^{2}&~{}&\cdots&\lambda^{d-1}\\ 0&\binom{1}{1}&\binom{2}{1}\lambda&~{}&\cdots&\binom{d}{1}\lambda^{d-2}\\ 0&0&\binom{2}{2}&~{}&\cdots&\binom{d}{2}\lambda^{d-3}\\ ~{}&~{}&~{}&~{}&\vdots&~{}\\ 0&\cdots&0&\binom{m-1}{m-1}&\cdots&\binom{d}{m-1}\lambda^{d-m}\end{bmatrix}.$
Let $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$ be a multiset consisting of distinct numbers $\lambda_{1},\dots,\lambda_{q}$ with multiplicities $m_{1},\dots,m_{q}$ such that $m_{1}+\cdots+m_{q}=d$. Define
$V_{\Lambda}^{(d)}=\begin{bmatrix}-&V_{\lambda_{1}^{(m_{1})}}^{(d)}&-\\ &\vdots&\\ -&V_{\lambda_{q}^{(m_{q})}}^{(d)}&-\\ \end{bmatrix}.$
The square matrix $V_{\Lambda}=V_{\Lambda}^{(d)}$ is called the confluent Vandermonde matrix (a.k.a. the generalized Vandermonde matrix) of $\Lambda$.
Example 2.2.
Suppose $p(x)=c_{0}+c_{1}x+\cdots+c_{4}x^{4}$ and $\Lambda=\{\lambda_{1}^{(1)},\lambda_{2}^{(2)},\lambda_{3}^{(2)}\}$. Then
$V_{\Lambda}=\begin{bmatrix}1&\lambda_{1}&\lambda_{1}^{2}&\lambda_{1}^{3}&\lambda_{1}^{4}\\ 1&\lambda_{2}&\lambda_{2}^{2}&\lambda_{2}^{3}&\lambda_{2}^{4}\\ 0&1&2\lambda_{2}&3\lambda_{2}^{2}&4\lambda_{2}^{3}\\ 1&\lambda_{3}&\lambda_{3}^{2}&\lambda_{3}^{3}&\lambda_{3}^{4}\\ 0&1&2\lambda_{3}&3\lambda_{3}^{2}&4\lambda_{3}^{3}\\ \end{bmatrix}.$
The confluent Vandermonde matrix $V_{\Lambda}$ of $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$ with $m_{1}+\cdots+m_{q}=d$ has the property that
$V_{\Lambda}\begin{bmatrix}c_{0}\\ \vdots\\ c_{d-1}\end{bmatrix}=\begin{bmatrix}\mathcal{D}^{(0)}p(\lambda_{1})\\ \vdots\\ \mathcal{D}^{(m_{1}-1)}p(\lambda_{1})\\ \vdots\\ \mathcal{D}^{(0)}p(\lambda_{q})\\ \vdots\\ \mathcal{D}^{(m_{q}-1)}p(\lambda_{q})\end{bmatrix}.$ (5)
Let $[p(x)]=[c_{0},\dots,c_{d-1}]^{\top}$ be the vector of coefficients of $p(x)=c_{0}+c_{1}x+\cdots+c_{d-1}x^{d-1}$ and $\operatorname{eval}_{\Lambda}(p(x))$ the the evaluations of $p(x)$ as shown in the right-hand side of Equation 5. Hence, we have $V_{\Lambda}[p(x)]=\operatorname{eval}_{\Lambda}(p(x))$ extending the classical result Equation 3.
In summary, $V_{\Lambda}$ is a matrix that transforms a polynomial into its evaluations at designated points and their derivatives according to the multiplicities.
3 Hasse derivative and confluent Vandermonde matrices over arbitrary fields
Let ${\mathbb{F}}[x]$ be the set of polynomials over a general field $\mathbb{F}$, and let $p(x)=c_{0}+c_{1}x+\cdots+c_{d-1}x^{d}\in{\mathbb{F}}[x]$. For any nonnegative integer $r$, define the Hasse derivative of $p(x)$ by
$\mathcal{D}^{(r)}p(x)=\sum_{i=r}^{d-1}c_{i}\binom{i}{r}x^{i-r}.$
When $\mathbb{F}=\mathbb{R}$, it is the same as the definition $\mathcal{D}^{(r)}=\frac{1}{r!}\cdot\frac{d^{r}}{dx^{r}}$ mentioned in the last section.
Note that the binomial coefficient $\binom{n}{k}$ is the coefficient of $x^{k}$ in the expansion of $(1+x)^{n}$, which is well defined in any field and is the same as $\frac{n!}{k!(n-k)!}$ when $\mathbb{F}$ is of characteristic 00. For this sake, an alternative definition of the Hasse derivative is $\mathcal{D}^{(r)}p(x)=[y^{r}]p(x+y)$, which is the coefficient of $y^{r}$ in $p(x+y)$ when treated as a polynomial of $y$. This can be verified by direct computation as follows.
$[y^{r}]p(x+y)$ $=[y^{r}]\sum_{i=0}^{d-1}c_{i}(x+y)^{i}=\sum_{i=0}^{d-1}c_{i}\cdot[y^{r}](x+y)^{i}$
$=\sum_{i=r}^{d-1}c_{i}\cdot[y^{r}](x+y)^{i}=\sum_{i=r}^{d-1}c_{i}\cdot\binom{i}{r}x^{i-r}=\mathcal{D}^{(r)}p(x)$
More intuitively, the evaluation $\mathcal{D}^{(r)}p(\lambda)$ is exactly the coefficient of $(x-\lambda)^{r}$ in the “Taylor’s expansion” of $p(x)$ at $\lambda$ as shown in the following.
Proposition 3.1.
Let $\lambda\in\mathbb{F}$. When $p(x)=c_{0}+c_{1}x+\cdots+c_{d-1}x^{d-1}$ is written as
$p(x)=a_{0}+a_{1}(x-\lambda)+\cdots+a_{d-1}(x-\lambda)^{d-1}.$
Then
$\mathcal{D}^{(r)}p(x)=\sum_{i=r}^{d-1}a_{i}\binom{i}{r}(x-\lambda)^{i-r}.$
Therefore, $\mathcal{D}^{(r)}p(\lambda)=a_{r}$ can be viewed as the $r$-th order term of $p(x)$ in its Taylor’s expansion at $\lambda$.
Proof.
With the alternative definition, we may compute
$\mathcal{D}^{(r)}p(x)$ $=[y^{r}]p(x+y)=[y^{r}]\sum_{i=0}^{d-1}a_{i}(x+y-\lambda)^{i}$
$=\sum_{i=0}^{d-1}a_{i}[y^{r}](x-\lambda+y)^{i}=\sum_{i=r}^{d-1}a_{i}[y^{r}](x-\lambda+y)^{i}$
$=\sum_{i=r}^{d-1}a_{i}\binom{i}{r}(x-\lambda)^{i-r}.$
When $\lambda$ is plugged into $\mathcal{D}^{(r)}p(x)$, each summand in the formula is zero except for $a_{r}\binom{r}{r}(x-\lambda)^{r-r}=a_{r}$. ∎
This view of Taylor’s expansion leads to the following identities.
Corollary 3.2.
Let $p(x)$ be a polynomial. Then
$\mathcal{D}^{(r)}(x-\lambda)^{m}p(x)\big{|}_{x=\lambda}=\begin{cases}\mathcal{D}^{(r-m)}p(\lambda)&\text{ if }r\geq m,\\ 0&\text{ otherwise.}\end{cases}$
In the subsequent discussion, $\mathcal{D}^{(r)}$ will always be treated as the Hasse derivative so that results obtained are valid for arbitrary fields. The simple observation in the corollary turns out to be a very useful tool for proving results on the confluent Vandermonde matrix as we shall see.
One may see [9] for more information about the Hasse derivative.
4 LU factorization
In the subsequent discussion, we shall always let $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$ for $q$ distinct elements in $\lambda_{1},\dots,\lambda_{q}$ in a field ${\mathbb{F}}$ with multiplicities $m_{1},\dots,m_{q}$ such that $d=m_{1}+\cdots+m_{q}$. We will find an LU factorization of the confluent Vandermonde matrix $V_{\Lambda}$ in the following. To achieve our goal, consider the $d+1$ polynomials:
$p_{0}(x)=1,\ p_{1}(x)=(x-\lambda_{1}),\ \ldots,\ p_{m_{1}}(x)=(x-\lambda_{1})^{m_{1}},$
$p_{m_{1}+1}=(x-\lambda_{1})^{m_{1}}(x-\lambda_{2}),\ \dots,\ p_{d}(x)=(x-\lambda_{1})^{m_{1}}\cdots(x-\lambda_{q})^{m_{q}}.$
Thus, $p_{i}(x)$ is a degree $i$ polynomial with leading coefficient equal to 1. Here we only use $p_{0}(x),\dots,p_{d-1}(x)$, and consider
$V_{\Lambda}\begin{bmatrix}|&~{}&|\\ [p_{0}(x)]&\cdots&[p_{d-1}(x)]\\ |&~{}&|\\ \end{bmatrix}=\begin{bmatrix}|&~{}&|\\ \operatorname{eval}_{\Lambda}(p_{0}(x))&\cdots&\operatorname{eval}_{\Lambda}(p_{d-1}(x))\\ |&~{}&|\\ \end{bmatrix}.$ (6)
Let $U^{-1}=\Big{[}[p_{0}(x)]\ \cdots\ [p_{d-1}(x)]\Big{]}$ and $L=\Big{[}\operatorname{eval}_{\Lambda}(p_{0}(x))\ \cdots\ \operatorname{eval}_{\Lambda}(p_{d-1}(x))\Big{]}$. Note that $U^{-1}$ is an upper triangular matrix with $1$’s on the diagonal. Thus, we can write it as the inverse of some matrix $U$, where $U=(U^{-1})^{-1}$ is an upper triangular matrix. Moreover, by Corollary 3.2, $L$ is a lower triangular matrix with the diagonal entries $1$ for $m_{1}$ times, $(\lambda_{2}-\lambda_{1})^{m_{1}}$ for $m_{2}$ times, $(\lambda_{3}-\lambda_{1})^{m_{1}}(\lambda_{3}-\lambda_{2})^{m_{2}}$ for $m_{3}$ times, and so on. Therefore, we have the following.
Theorem 4.1.
Equation 6, written as $V_{\Lambda}U^{-1}=L$, leads to the LU factorization $V_{\Lambda}=LU$.
Theorem 4.2.
For any $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$, $\det(V_{\Lambda})=\prod_{\begin{subarray}{c}i,j\\ i<j\end{subarray}}(\lambda_{j}-\lambda_{i})^{m_{i}m_{j}}$. Consequently, $V_{\Lambda}$ is always invertible.
Proof.
This is immediate from the fact that $V_{\Lambda}U^{-1}=L$ and that the determinant of an upper triangular matrix (or a lower triangular matrix) is the product of its diagonal entries. ∎
Using Example 2.2 with $(\lambda_{1},\lambda_{2},\lambda_{3})=(1,3,6)$, we have the following numerical example.
Example 4.3.
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$. Then
$p_{0}(x)=1,\ p_{1}(x)=(x-1),\ p_{2}(x)=(x-1)(x-3),$
$p_{3}(x)=(x-1)(x-3)^{2},\ p_{4}(x)=(x-1)(x-3)^{2}(x-6).$
Also,
$V_{\Lambda}=\begin{bmatrix}1&1&1&1&1\\ 1&3&9&27&81\\ 0&1&6&27&108\\ 1&6&36&216&1296\\ 0&1&12&108&864\end{bmatrix}\text{ and }U^{-1}=\begin{bmatrix}1&-1&3&-9&54\\ 0&1&-4&15&-99\\ 0&0&1&-7&57\\ 0&0&0&1&-13\\ 0&0&0&0&1\end{bmatrix}.$
Thus,
$V_{\Lambda}U^{-1}=\begin{bmatrix}p_{0}(1)&\cdots&p_{4}(1)\\ p_{0}(3)&\cdots&p_{4}(3)\\ \mathcal{D}^{(1)}p_{0}(3)&\cdots&\mathcal{D}^{(1)}p_{4}(3)\\ p_{0}(6)&\cdots&p_{4}(6)\\ \mathcal{D}^{(1)}p_{0}(6)&\cdots&\mathcal{D}^{(1)}p_{4}(6)\\ \end{bmatrix}=\begin{bmatrix}1&0&0&0&0\\ 1&2&0&0&0\\ 0&1&2&0&0\\ 1&5&15&45&0\\ 0&1&8&39&45\end{bmatrix}=L.$
Since $\det(U^{-1})=1$, it follows that $\det(V_{\Lambda})=\det(L)=(3-1)^{2}(6-1)^{2}(6-3)^{4}$.
5 The Chinese Remainder Theorem
Let $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$ with $d=m_{1}+\cdots+m_{q}$. Define
$\hat{h}_{j,m}(x)=(x-\lambda_{j})^{m}\prod_{\begin{subarray}{c}k=1,\ldots,q\\ k\neq j\end{subarray}}(x-\lambda_{k})^{m_{k}}$
for $j=1,\ldots,q$ and $m=0,\ldots,m_{j}-1$. Note that there are $d$ polynomials. When we write $\hat{h}_{1,0}(x)$, $\ldots$, $\hat{h}_{q,m_{q}-1}(x)$, we mean $\hat{h}_{1,0}(x)$, $\ldots$, $\hat{h}_{1,m_{1}-1}(x)$, $\ldots$, $\hat{h}_{q,0}(x)$, $\ldots$, $\hat{h}_{q,m_{q}-1}(x)$ in order. The Chinese Remainder Theorem for polynomial rings asserts that every polynomial of degree less than $d$ can be uniquely written as a linear combination of ^$\hat{\mathcal{H}}_{\Lambda}=\{\hat{h}_{1,0}(x),\ldots,\hat{h}_{q,m_{q}-1}(x)\}$; equivalently, ^$\hat{\mathcal{H}}_{\Lambda}$ is a basis of the space of all polynomial of degree less than $d$. For example, using Example 2.2, we have the following.
Example 5.1.
Let $\Lambda=\{\lambda_{1}^{(1)},\lambda_{2}^{(2)},\lambda_{3}^{(2)}\}$. Then
$\hat{h}_{1,0}(x)=(x-\lambda_{2})^{2}(x-\lambda_{3})^{2},\hskip 1.5pt\in\ ~{}~{}~{}$
$\hat{h}_{2,0}(x)=(x-\lambda_{1})(x-\lambda_{3})^{2},\ \hat{h}_{2,1}(x)=(x-\lambda_{2})(x-\lambda_{1})(x-\lambda_{3})^{2},$
$\hat{h}_{3,0}(x)=(x-\lambda_{1})(x-\lambda_{2})^{2},\ \hat{h}_{3,1}(x)=(x-\lambda_{3})(x-\lambda_{1})(x-\lambda_{2})^{2}$
form a basis for the linear space of polynomial of degree less than 5.
The confluent Vandermonde matrix provides a simple way to prove this version of the Chinese Remainder Theorem as follows. Let
$\hat{H}_{\Lambda}=\begin{bmatrix}|&~{}&|\\ [\hat{h}_{1,0}]&\cdots&[\hat{h}_{q,m_{q}-1}]\\ |&~{}&|\\ \end{bmatrix}.$
Then the Chinese Remainder Theorem is equivalent to that $\hat{H}$ is invertibile. To see this, we observe that
$\mathcal{D}^{(r)}\hat{h}_{j,m}(\lambda_{i})=\begin{cases}0&\text{if }i\neq j\text{ and }r<m_{i},\\ 0&\text{if }i=j\text{ and }r<m,\\ \star&\text{if }i=j\text{ and }r=m,\\ ?&\text{if }i=j\text{ and }r>m.\\ \end{cases}$
Here $\star$ stands for a nonzero number and ???? stands for any value in the field. To be more precise, by Corollary 3.2 we have
$\mathcal{D}^{(m)}\hat{h}_{j,m}(\lambda_{j})=\mathcal{D}^{(0)}\prod_{\begin{subarray}{c}k=1,\ldots,q\\ k\neq j\end{subarray}}(x-\lambda_{k})^{m_{k}}\Big{|}_{x=\lambda_{j}}=\prod_{\begin{subarray}{c}k=1,\ldots,q\\ k\neq j\end{subarray}}(\lambda_{j}-\lambda_{k})^{m_{k}}\neq 0$
if $\lambda_{i}-\lambda_{j}\neq 0$ whenever $i\neq j$. Consequently, $V_{\Lambda}\hat{H}$ is a lower triangular matrix with diagonal entries all nonzero. Therefore, both $V_{\Lambda}$ and $\hat{H}$ are invertible, proving the Chinese Remainder Theorem.
Note that this argument also proves the invertibility condition of $V_{\Lambda}$ without using the formula in Theorem 4.2.
Again, using Example 2.2 with $(\lambda_{1},\lambda_{2},\lambda_{3})=(1,3,6)$, we have the following.
Example 5.2.
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$. Then
$\hat{h}_{1,0}(x)$ $=(x-1)^{0}(x-3)^{2}(x-6)^{2},$
$\hat{h}_{2,0}(x)$ $=(x-1)(x-3)^{0}(x-6)^{2},$
$\hat{h}_{2,1}(x)$ $=(x-1)(x-3)^{1}(x-6)^{2},$
$\hat{h}_{3,0}(x)$ $=(x-1)(x-3)^{2}(x-6)^{0},$
$\hat{h}_{3,1}(x)$ $=(x-1)(x-3)^{2}(x-6)^{1}.$
We may compute that
$(x-1)(x-6)^{2}$ $=(x-3+2)(x-3-3)^{2}$
$=18-3(x-3)-4(x-3)^{2}+(x-3)^{3},$
so $\mathcal{D}^{(0)}\hat{h}_{2,0}(3)=18$, $\mathcal{D}^{(1)}\hat{h}_{2,0}(3)=-3$, and $\mathcal{D}^{(0)}\hat{h}_{2,1}(3)=0$, $\mathcal{D}^{(1)}\hat{h}_{2,1}(3)=18$. Other evaluations can be done in similar manners. Therefore,
$V_{\Lambda}=\begin{bmatrix}1&1&1&1&1\\ 1&3&9&27&81\\ 0&1&6&27&108\\ 1&6&36&216&1296\\ 0&1&12&108&864\end{bmatrix}\text{ and }\hat{H}_{\Lambda}=\begin{bmatrix}324&-36&108&-9&54\\ -324&48&-180&15&-99\\ 117&-13&87&-7&57\\ -18&1&-16&1&-13\\ 1&0&1&0&1\end{bmatrix}.$
Also,
$V_{\Lambda}\hat{H}_{\Lambda}=\begin{bmatrix}\hat{h}_{1,0}(1)&\cdots&\hat{h}_{3,1}(1)\\ \hat{h}_{1,0}(3)&\cdots&\hat{h}_{3,1}(3)\\ \mathcal{D}^{(1)}\hat{h}_{1,0}(3)&\cdots&\mathcal{D}^{(1)}\hat{h}_{3,1}(3)\\ \hat{h}_{1,0}(6)&\cdots&\hat{h}_{3,1}(6)\\ \mathcal{D}^{(1)}\hat{h}_{1,0}(6)&\cdots&\mathcal{D}^{(1)}\hat{h}_{3,1}(6)\\ \end{bmatrix}=\begin{bmatrix}100&0&0&0&0\\ 0&18&0&0&0\\ 0&-3&18&0&0\\ 0&0&0&45&0\\ 0&0&0&39&45\end{bmatrix}.$
Theorem 5.3.
Let $\Lambda=\{\lambda_{1}^{(m_{1})},\dots,\lambda_{q}^{(m_{q})}\}$ with $m_{1}+\cdots+m_{q}=d$. Then $T_{\Lambda}=V_{\Lambda}\hat{H}_{\Lambda}$ is lower triangular matrix and is a block-diagonal matrix with $q$ blocks. For any given polynomial $p(x)$ of degree less than $d$, $\hat{H}^{-1}_{\Lambda}[p(x)]=T^{-1}_{\Lambda}V_{\Lambda}[p(x)]$ records the coefficients of $p(x)$ as a linear combination of ^$\hat{\mathcal{H}}_{\Lambda}$.
6 Hermite polynomials
Let $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$ with $d=m_{1}+\cdots+m_{q}$. Let $H_{\Lambda}=V_{\Lambda}^{-1}$. Then we may find polynomials $\mathcal{H}_{\Lambda}=\{h_{1,0}(x),\ldots,h_{q,m_{q}-1}(x)\}$ such that
$H_{\Lambda}=\begin{bmatrix}|&~{}&|\\ [h_{1,0}(x)]&\cdots&[h_{q,m_{q}-1}(x)]\\ |&~{}&|\\ \end{bmatrix}.$
The set of polynomials $\mathcal{H}_{\Lambda}$ is called the Hermite polynomials of $\Lambda$. Given any polynomial $p(x)$ of degree less than $d$, one may recover the polynomial from the information $\operatorname{eval}_{\Lambda}(p(x))$ by
$[p(x)]=V_{\Lambda}^{-1}\operatorname{eval}_{\Lambda}(p(x))=H_{\Lambda}\operatorname{eval}_{\Lambda}(p(x))$
and write $p(x)$ as a linear combination of polynomials in $\mathcal{H}_{\Lambda}$.
Example 6.1.
Let $\Lambda=\{1^{(1)},3^{(2)},6^{(2)}\}$. Suppose $p(x)\in{\mathbb{F}}[x]$ is a polynomial of degree less than $5$, and $p(1)=c_{1}$, $p(3)=c_{2}$, $\mathcal{D}^{(1)}(p(3))=c_{3}$, $p(6)=c_{4}$, $\mathcal{D}^{(1)}(p(6))=c_{5}$ are given. Then
$V_{\Lambda}[p(x)]=\operatorname{eval}_{\Lambda}(p(x))=\begin{bmatrix}c_{1}\\ c_{2}\\ c_{3}\\ c_{4}\\ c_{5}\end{bmatrix}.$
With $H_{\Lambda}=V_{\Lambda}^{-1}$, we have $H\operatorname{eval}_{\Lambda}(p(x))=[p(x)]$. Moreover, by Example 5.2, we may compute $T_{\Lambda}=V_{\Lambda}\hat{H}_{\Lambda}$ and
$T^{-1}_{\Lambda}=\begin{bmatrix}\frac{1}{100}&0&0&0&0\\ 0&\frac{1}{18}&0&0&0\\ 0&\frac{1}{108}&\frac{1}{18}&0&0\\ 0&0&0&\frac{1}{45}&0\\ 0&0&0&-\frac{13}{675}&\frac{1}{45}\end{bmatrix}$
Let $\hat{h}_{1,0},\ldots,\hat{h}_{3,1}$ be as in Example 5.2. Since $H_{\Lambda}=V_{\Lambda}^{-1}=\hat{H}_{\Lambda}T^{-1}_{\Lambda}$, the Hermite polynomials are
$h_{1,0}(x)$ $=\frac{1}{100}\hat{h}_{1,0}(x)=\left[\frac{1}{100}\right](x-3)^{2}(x-6)^{2},$
$h_{2,0}(x)$ $=\frac{1}{18}\hat{h}_{2,0}(x)+\frac{1}{108}h_{2,1}(x)=(x-1)\left[\frac{1}{18}+\frac{1}{108}(x-3)\right](x-6)^{2},$
$h_{2,1}(x)$ $=\frac{1}{18}\hat{h}_{2,1}(x)=(x-1)\left[\frac{1}{18}(x-3)\right](x-6)^{2},$
$h_{3,0}(x)$ $=\frac{1}{45}\hat{h}_{3,0}(x)-\frac{13}{675}\hat{h}_{3,1}(x)=(x-1)(x-3)^{2}\left[\frac{1}{45}-\frac{13}{675}(x-6)\right],$
$h_{3,1}(x)$ $=\frac{1}{45}\hat{h}_{3,1}(x)=(x-1)(x-3)^{2}\left[\frac{1}{45}(x-6)\right].$
For each of these polynomials, its evaluations given by $\Lambda$ are all 00 except of a unique $1$. Therefore, for any given evaluations $c_{1},\ldots,c_{5}$, we have the Hermite interpolation
$p(x)=c_{1}h_{1,0}(x)+c_{2}h_{2,0}(x)+c_{3}h_{2,1}(x)+c_{4}h_{3,0}(x)+c_{5}h_{3,1}(x).$
Note that computing $T^{-1}_{\Lambda}$ is much easier than computing $V_{\Lambda}^{-1}$. In fact, the author of [13] used similar ideas to compute the inverse of a confluent Vandermonde matrix. The Hermite interpolation is also used to calculate the functions of companion matrices [5, 16].
7 Companion matrix
Let
$p(x)=c_{0}+c_{1}x+\cdots+c_{d}x^{d}\text{ with }c_{d}=1$
be a polynomial of degree $d$. The companion matrix of $p(x)$ is defined as
$C=\begin{bmatrix}0&\cdots&0&-c_{0}\\ 1&\ddots&\vdots&-c_{1}\\ ~{}&\ddots&0&\vdots\\ 0&~{}&1&-c_{d-1}\end{bmatrix}$
When
$p(x)=\prod_{\lambda\in\Lambda}(x-\lambda).$
is the monic polynomial with roots in $\Lambda$, we write $C_{\Lambda}$ as the companion matrix of $p(x)$. Here, we always assume that the roots lie in the field. In particular, the discussion always work for an algebraically closed field.
Suppose $\Lambda=\{\lambda_{1}^{(m_{1})},\ldots,\lambda_{q}^{(m_{q})}\}$. Let $J_{\lambda,m}$ be the $m\times m$ lower-triangular Jordan block of $\lambda$ and define
$J_{\Lambda}=\begin{bmatrix}J_{\lambda_{1},m_{1}}&~{}&~{}\\ ~{}&\ddots&~{}\\ ~{}&~{}&J_{\lambda_{q},m_{q}}\end{bmatrix}.$
Brand[1] stated that
$V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}=J_{\Lambda}$
is the Jordan canonical form of $C_{\Lambda}$ with exactly $1$ Jordan block for each eigenvalue. In particular, when all elements in $\Lambda$ are distinct, $V_{\Lambda}C_{\Lambda}V_{\lambda}^{-1}$ is a diagonal matrix.
Example 7.1.
Let $f(x)=(x-\lambda_{1})(x-\lambda_{2})^{2}(x-\lambda_{3})^{2}$ such that $\lambda_{i}-\lambda_{j}\neq 0$ whenever $i\neq j$. Then $V_{\Lambda}C_{f}V_{\lambda}^{-1}$ will be a direct sum of $[\lambda_{1}],\begin{bmatrix}\lambda_{2}&0\cr 1&\lambda_{2}\end{bmatrix}$ and $\begin{bmatrix}\lambda_{3}&0\cr 1&\lambda_{3}\end{bmatrix}$.
This Jordan form result is elegant, but no proof was given in [1]. With the discussions in the previous sections, we can prove the Jordan form result above and gain better insight into it. Recall that $\operatorname{eval}_{\Lambda}(f(x))$ is a vector evaluating the polynomial $f(x)$. By definition, we have $\operatorname{eval}_{\Lambda}(p(x))=\mathbf{0}$. Moreover, each evaluation has the following properties.
Proposition 7.2.
Let $f(x)\in{\mathbb{F}}[x]$ with distinct roots $\lambda_{1},\dots,\lambda_{q}$. Then
  1. 1.
    $\mathcal{D}^{(0)}(xf(x))\big{|}_{x=\lambda}=\lambda\mathcal{D}^{(0)}(f(x))\big{|}_{x=\lambda}$,
  2. 2.
    $\mathcal{D}^{(r)}(xf(x))\big{|}_{x=\lambda}=\mathcal{D}^{(r-1)}(f(x))\big{|}_{x=\lambda}+\lambda\mathcal{D}^{(r)}(f(x))\big{|}_{x=\lambda}$ if $r\geq 1$, and
  3. 3.
    $\operatorname{eval}_{\Lambda}(xf(x))=J_{\Lambda}\operatorname{eval}_{\Lambda}(f(x))$.
Proof.
The first identity holds since both sides equal to $\lambda f(\lambda)$.
For the field of real numbers, the second identity could be written as
$\frac{1}{r!}\cdot\frac{d^{r}}{dx^{r}}(xf(x))\Big{|}_{x=\lambda}=\frac{1}{r!}\left[\binom{r}{1}\frac{d^{r-1}}{dx^{r-1}}(f(x))+x\cdot\frac{d^{r}}{dx^{r}}(xf(x))\right]\Big{|}_{x=\lambda}$
and is the product rule applied to $xf(x)$. However, we emphasize that the identity remains valid from the Hasse derivative point of view, and the proof is even more intuitive. By direct computation and Corollary 3.2, we have
$\mathcal{D}^{(r)}(xf(x))\big{|}_{x=\lambda}$ $=\mathcal{D}^{(r)}((x-\lambda)f(x)+\lambda f(x))\big{|}_{x=\lambda}$
$=\mathcal{D}^{(r)}((x-\lambda)f(x))\big{|}_{x=\lambda}+\mathcal{D}^{(r)}(\lambda f(x))\big{|}_{x=\lambda}$
$=\mathcal{D}^{(r-1)}(f(x))\big{|}_{x=\lambda}+\lambda\mathcal{D}^{(r)}(f(x))\big{|}_{x=\lambda}$
Finally, the third identity is simply a matrix-vector form of the first two identities. ∎
Let $p_{j}(x)=x^{j}$. Then the columns of the $d\times d$ identity matrix $I$ are $[p_{0}(x)],\ldots,[p_{d-1}(x)]$ and
$C_{\Lambda}=\begin{bmatrix}|&~{}&|&|\\ [xp_{0}(x)]&\cdots&[xp_{d-2}(x)]&[xp_{d-1}(x)-p(x)]\\ |&~{}&|&|\\ \end{bmatrix}.$
With Proposition 7.2 and the fact $\operatorname{eval}_{\Lambda}(p(x))=\mathbf{0}$, we have
$V_{\Lambda}C_{\Lambda}=J_{\Lambda}\begin{bmatrix}|&~{}&|\\ \operatorname{eval}_{\Lambda}(p_{0}(x))&~{}&\operatorname{eval}_{\Lambda}(p_{d-1}(x))\\ |&~{}&|\\ \end{bmatrix}=J_{\Lambda}V_{\Lambda}I=J_{\Lambda}V_{\Lambda}.$
Therefore, $V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}=J_{\Lambda}$ follows.
Moreover, with our discussion in earlier sections, $V_{\Lambda}^{-1}$ comes from the Lagrange polynomials when all elements in $\Lambda^{-1}$ are distinct and from the Hermite polynomials in general. Let $\mathcal{E}_{d}=\{p_{0}(x),\ldots,p_{d-1}(x)\}$ and $\mathcal{H}_{\Lambda}$ the Hermite polynomials. Then both $\mathcal{E}_{d}$ and $\mathcal{H}_{\Lambda}$ are bases of the space $\mathcal{P}_{d-1}$ of polynomials of degree less than $d$. Consider the linear transformation $\phi:\mathcal{P}_{d-1}\rightarrow\mathcal{P}_{d-1}$ defined by $\phi(f(x))=xf(x)\pmod{p(x)}$. Then $C_{\Lambda}$ can be viewed as the matrix representation of $\phi$ with respect to $\mathcal{E}_{d}$, while $J_{\Lambda}$ is the matrix representation of $\phi$ with respect to $\mathcal{H}_{\Lambda}$. Thus, $V_{\Lambda}$ and $H_{\Lambda}=V_{\Lambda}^{-1}$ are the change of basis matrices between $\mathcal{E}_{d}$ and $\mathcal{H}_{\Lambda}$. This provides a comprehensive understanding of the Jordan decomposition of $C_{\Lambda}$.
The decomposition $V_{\Lambda}C_{\Lambda}V_{\Lambda}^{-1}=J_{\Lambda}$ was then used for solving differential equations in [2] and yet more applications in [16, 21, 4].
Moreover, Gilbert [8] and the references therein considered integer matrices with integer eigenvalues with integer (left and right) generalized eigenvectors. The paper pointed out that $C_{\Lambda}$ is one of such matrices when $\Lambda$ contains only two distinct values that are integers with difference $1$. This can again be seen from our discussion. When $\Lambda$ contains only one integer value or two distinct integer values with difference $1$, we have $\det(V_{\Lambda})=1$ and $V_{\Lambda}^{-1}$ is an integer matrix. In these cases, $V_{\Lambda}$, $C_{\Lambda}$, $V_{\Lambda}^{-1}$, and $J_{\Lambda}$ are all integer matrices.